A cube made of an unknown material has a dimension of 6 cm on all edges. If the upper face is displaced laterally by 0.1 mm by a force of 2 x 106 dynes acting parallel to the surface, calculate the shear modulus of the material


1
Expert's answer
2022-03-09T10:32:44-0500

The shear modulus of the material is given by

"G=\frac{Fl}{A\Delta x}"

"G=\rm\frac{20\:N*0.06\: m}{(0.06\: m)^2*0.1*10^{-3}\: m}=3.3*10^{6}\:\rm Pa"